Grade 12th passPhysical Chemistry

How many cc of oxygen will be liberated by 2 ampere current flowing for 3 minutes and 13 seconds through acidulated water

Nimra
8 Years agoGrade 12th pass

1 Answer

Arun
8 Years ago
Dear student
 
Q = 2 * 193 C = 386 C
 
H2O ---------> H2 + ½ O2
 
O2 – -------> ½ O2 + 2 e-
 
2F = 2* 96500 C gives
O2 = ½ mole = 11200 cc
 
hence
 
386 C will give O2 = 11200 * 386 / (2 * 96500) = 22.4 cc
 
Regards
Arun (askIITians forum expert)